If A and B are the roots of the equation x^2+Ax+B=0 then find the minimum value of this equation.

Grade:11

1 Answers

Aman Dodabetta
7 years ago
The minimum value of the equation is given by -D/4a.
D = A^2-4AB, so -D = 4AB – A^2
4a = 4, so -D/a=4a = 4AB-A^2/4.
Since A and B are the roots, 4AB-A^2 = 0
So minimum value of equation is equal to 0